OpenStudy (anonymous):

how do you find the inverse of f(x): f(x)=2/3x+6

OpenStudy (anonymous):

It's a 4 step process

OpenStudy (anonymous):

First, replace f(x) with "y"

OpenStudy (anonymous):

\[f(x)=\frac{2}{3}x+6\] or \[f(x)=\frac{2}{3x+6}\]

OpenStudy (anonymous):

So you've got f(x) = 2/3x+6 y = 2/3x + 6 Next, the "x" and "y" switch places.

OpenStudy (anonymous):

So you've got: x = 2/3y +6 Next, you solve for "y"

OpenStudy (anonymous):

x=2/3y+6 Subtract 6 from both sides x-6 = 2/3y Then, you can multiply by 3/2 on both sides, and you'll have that y by itself. 3/2(x-6) = y

OpenStudy (anonymous):

And the last step is to replace y with: \[ f^{-1}(x)\]

OpenStudy (anonymous):

So you're left with: \[f^{-1}(x)=\frac{3(x-6)}{2}\]

OpenStudy (anonymous):

Or, if you'd prefer: \[ f^{-1}(x)=\frac{3}{2}x - 9\]